# Commutation relations for two-sided radial SLE

### v2

@@ 3.3 Resampling property of two-sided radial SLE with spiral
The image $\tilde{U}_{t_{1}}:=g_{t_{1}}^{(1)}(U_{t_{1}})$ is bounded by $I_{t_{1}}$ and a circular arc $\alpha_{t_{1}}$ meeting the endpoints of $I_{t_{1}}$ with angle $\lambda/\pi$. Since $|I^{c}_{t_{1}}|\to 0$, the diameter of the domain $\mathbb{D}\setminus\tilde{U}_{t_{1}}$, which contains $\tilde{\eta}^{(2)}_{[0,t_{2}]}=g_{t_{1}}^{(1)}\left(\eta^{(2)}_{[0,t_{2}]}\right)$, converges to 00.
@@
Recall that the map $g_{\mathbf{t},1}$ maps out the curve $\tilde{\eta}^{(2)}_{[0,t_{2}]}=g_{t_{1}}^{(1)}\left(\eta_{[0,t_{2}]}^{(2)}\right)$. If we conjugate $g_{\mathbf{t},1}$ by the rotation $\rho_{t_{1}}:\mathbb{D}\to\mathbb{D}$ such that the image of the mid-point of $\alpha_{t_{1}}$ under $\rho_{t_{1}}$ lies in $(0,1)$ (so that $R(\alpha_{t_{1}})$ is symmetric with respect to the real line and $\rho_{t_{1}}(\alpha_{t_{1}})$ shrinks to the point $1\in\partial\mathbb{D}$), the map $\tilde{g}_{\mathbf{t},1}:=\rho_{t_{1}}\circ g_{\mathbf{t},1}\circ\rho_{t_{1}}^{-1}$ converges in Carathéodory topology (namely, uniformly on compact subsets) to the identity map in $\mathbb{D}$ as $t_{1}\to\infty$. If we Schwarz-reflect $g_{\mathbf{t},1}$ along $\partial\mathbb{D}\setminus g_{t_{1}}^{(1)}(\exp(\mathfrak{i}\theta_{2}))$, we see that the convergence also extends to the boundary, more precisely, we obtain that $\tilde{g}_{\mathbf{t},1}$ converges uniformly on all compact subsets of $\overline{\mathbb{D}}\setminus\{1\}$ (and the map is well-defined on every such compact subset for large enough $t_{1}$), so do the derivatives of $\tilde{g}_{\mathbf{t},1}$ with respect to $z$.
Finally, since $\Delta_{t_{1}}$ is bounded away from 00 and $2\pi$ almost surely, as we proved above, we obtain that $R_{\mathbf{t}}\to 1$ almost surely, which completes the proof. ∎
@@ 3.4 Chordal SLE weighted by conformal radius
In this section, we show that the partition functions $\mathcal{Z}_{\alpha}$ correspond to the chordal SLE weighted by the conformal radius to the power $-\alpha$.
For this, we first calculate the Laplace transform of the conformal radius of the complement of chordal SLE. Usually, chordal SLE is defined in the upper-half plane as in Appendix B. It is more convenient here to describe it in the unit disc $\mathbb{D}$ via a change of coordinate. Fix $\kappa\in(0,8)$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. Suppose $\gamma$ is chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$. We parameterize it by the capacity and define $g_{t},\xi_{t}$ accordingly as in Section 3.1. Denote by $T$ the first time $\gamma$ disconnects $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ from the origin. A chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$, up to $T$, has the same law as radial $\operatorname{SLE}_{\kappa}(\kappa-6)$ starting from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{2}}$, up to the same time, see [36]. In other words, its driving function $\xi_{t}$ solves the following SDE:
@@
We first show that $\Phi(\kappa,\alpha;u)$ is finite as long as $\alpha<1-\kappa/8$. This is done in [31, Proof of Prop. 3.5]. For the reader’s convenience, we summarize its proof here.
@@
The disconnection time $T$ is the first time that $\theta_{t}$ hits 00 or $2\pi$. Suppose $f$ is an analytic function defined on $(0,1)$. Then $\mathrm{e}^{\alpha t}f\left(\left(\sin(\theta_{t}/4)\right)^{2}\right)$ is a local martingale if and only if $f$ satisfies (3.19). Since $f_{1},f_{2}$ are solutions to this ODE, the processes
@@ 3.5 Commuting SLEs without interchangeability
$\mathrm{CR}(\mathbb{D}\setminus\gamma)^{-\alpha}(\beta 1_{\{0\text{ is to the left of }\gamma\}}+(1-\beta)1_{\{0\text{ is to the right of }\gamma\}}),$
@@
$\mathrm{CR}(\mathbb{D}\setminus\eta^{(2)})^{-\alpha}(\beta 1_{\{0\text{ is to the right of }\eta^{(2)}\}}+(1-\beta)1_{\{0\text{ is to the left of }\eta^{(2)}\}}).$
@@ 4.1 Commutation relation when $\kappa=0$
We consider an interchangeable and locally commuting $2$-radial $\operatorname{SLE}_{0}$. Let $b_{1},b_{2}:S^{1}\times S^{1}\setminus\Delta\to\mathbb{R}$ be $C^{2}$ functions as in the condition (MARG). Then (2.4) and (2.6) imply that there exists $\mathcal{U}:\{(\theta_{1},\theta_{2})\in\mathbb{R}^{2}\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}$ and a constant $C$ such that
@@
$\begin{split}(\partial_{2}\mathcal{U})^{2}+2\cot(\theta_{12}/2)\partial_{1}\mathcal{U}-\frac{3}{\left(\sin(\theta_{12}/2)\right)^{2}}&=C,\\ (\partial_{1}\mathcal{U})^{2}+2\cot(\theta_{21}/2)\partial_{2}\mathcal{U}-\frac{3}{\left(\sin(\theta_{21}/2)\right)^{2}}&=C,\end{split}$ (4.1)
@@
The proof of Proposition 2.3 up to (2.11) holds verbatim when $\kappa=0$.
From (2.9) we know that there is a function $\mathcal{U}:\{(\theta_{1},\theta_{2})\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}$ such that we can write $b_{1}=\partial_{1}\mathcal{U}$ and $b_{2}=\partial_{2}\mathcal{U}$. Equations (2.10) and (2.11) give:
$\partial_{1}\left((\partial_{2}\mathcal{U})^{2}+2\cot(\theta_{12}/2)\partial_{1}\mathcal{U}-\frac{3}{\left(\sin(\theta_{12}/2)\right)^{2}}\right)=0$ (4.4)
$\partial_{2}\left((\partial_{1}\mathcal{U})^{2}+2\cot(\theta_{21}/2)\partial_{2}\mathcal{U}-\frac{3}{\left(\sin(\theta_{21}/2)\right)^{2}}\right)=0.$ (4.5)
Hence, there exist functions $C_{1}(\theta_{1})$ and $C_{2}(\theta_{2})$ such that
$\begin{split}&(\partial_{2}\mathcal{U})^{2}+2\cot(\theta_{12}/2)\partial_{1}\mathcal{U}-\frac{3}{\left(\sin(\theta_{12}/2)\right)^{2}}=C_{2}(\theta_{2})\\ &(\partial_{1}\mathcal{U})^{2}+2\cot(\theta_{21}/2)\partial_{2}\mathcal{U}-\frac{3}{\left(\sin(\theta_{21}/2)\right)^{2}}=C_{1}(\theta_{1}).\end{split}$ (4.6)
Since $b_{j}$ is invariant under the rotation $(\theta_{1},\theta_{2})\mapsto(\theta_{1}+a,\theta_{2}+a)$, so are the left-hand sides of (4.6). Thus, $C_{1}(\theta_{1})=C_{1}$ and $C_{2}(\theta_{2})=C_{2}$ are constant. The rotational invariance of $b_{j}$ also gives
$0=\partial_{1}(\partial_{1}\mathcal{U}+\partial_{2}\mathcal{U})=\partial_{2}(\partial_{1}\mathcal{U}+\partial_{2}\mathcal{U}).$
We let $\mu\in\mathbb{R}$ such that
$2\mu\equiv\partial_{1}\mathcal{U}+\partial_{2}\mathcal{U}.$
Taking the difference of the equations in (4.6), we obtain
$\left(\partial_{1}\mathcal{U}-\partial_{2}\mathcal{U}+2\cot(\theta_{21}/2)\right)2\mu=C_{1}-C_{2}$ (4.7)
Case 1: If $\mu\neq 0$, then
$\begin{cases}\partial_{1}\mathcal{U}=\frac{C_{1}-C_{2}}{4\mu}+\mu-\cot(\theta_{21}/2),\\ \partial_{2}\mathcal{U}=\frac{C_{2}-C_{1}}{4\mu}+\mu+\cot(\theta_{21}/2).\end{cases}$
Plugging this back into (4.6) shows that $C_{1}=C_{2}=:C$ and $C=\mu^{2}-3.$ Therefore, up to an additive constant, we have
$\mathcal{U}(\theta_{1},\theta_{2})=2\log\sin(\theta_{21}/2)+\mu(\theta_{1}+\theta_{2}).$
Case 2: If $\mu=\partial_{1}\mathcal{U}+\partial_{2}\mathcal{U}=0$, then it follows directly that $C_{1}=C_{2}=:C$. Moreover, $\mathcal{U}$ only depends on $\theta:=\theta_{2}-\theta_{1}$ and writing $\mathcal{U}(\theta)=\mathcal{U}(\theta_{1},\theta_{2})$ with a slight abuse of notation, we have from the interchangeability condition that
$\mathcal{U}(\theta)=\mathcal{U}(2\pi-\theta).$ (4.8)
Thus, (4.1) becomes
$(\mathcal{U}^{\prime}(\theta)+\cot(\theta/2))^{2}=C+\frac{3+\left(\cos(\theta/2)\right)^{2}}{\left(\sin(\theta/2)\right)^{2}}.$ (4.9)
Eq. (4.9) gives the solutions
$\mathcal{U}^{\prime}(\theta)=\pm\sqrt{C+\frac{3+\left(\cos(\theta/2)\right)^{2}}{\left(\sin(\theta/2)\right)^{2}}}-\cot(\theta/2).$ (4.10)
Since $\mathcal{U}^{\prime}(\pi)=0$ from (4.8), this implies that $C=-3$ and
$\mathcal{U}^{\prime}(\theta)=\cot(\theta/2)\text{ or }-3\cot(\theta/2).$
Namely,
$\mathcal{U}(\theta)=2\log\sin(\theta/2)\text{ or }-6\log\sin(\theta/2)$
up to an additive constant. ∎
@@
We note that when $\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{0}(\theta_{1},\theta_{2})=2\log\sin(\theta_{21}/2)$, the corresponding commuting $\operatorname{SLE}_{0}$ is the two-sided radial $\operatorname{SLE}_{0}$ (with zero spiraling rate). This is also the geodesic pair in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ studied in [38, 24, 25], or in other words, the concatenation of $\eta^{(1)}$ and $\eta^{(2)}$ is the chord minimizing the chordal Loewner energy among all chords in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ passing through 00, hence can be viewed as the chordal $\operatorname{SLE}_{0}$ “conditioned” to pass through 00.

### v3

@@ 3.3 Resampling property of two-sided radial SLE with spiral
The image $\tilde{U}_{t_{1}}:=g_{t_{1}}^{(1)}(U_{t_{1}})$ is bounded by $I_{t_{1}}$ and a circular arc $\alpha_{t_{1}}$ meeting the endpoints of $I_{t_{1}}$ with angle $\lambda/\pi$. Since $|I^{c}_{t_{1}}|\to 0$, the diameter of the domain $\mathbb{D}\setminus\tilde{U}_{t_{1}}$, which contains $\tilde{\eta}^{(2)}_{[0,t_{2}]}=g_{t_{1}}^{(1)}\left(\eta^{(2)}_{[0,t_{2}]}\right)$, converges to 0.
@@
Recall that the map $g_{\mathbf{t},1}$ maps out the curve $\tilde{\eta}^{(2)}_{[0,t_{2}]}=g_{t_{1}}^{(1)}\left(\eta_{[0,t_{2}]}^{(2)}\right)$. If we conjugate $g_{\mathbf{t},1}$ by the rotation $\rho_{t_{1}}:\mathbb{D}\to\mathbb{D}$ such that the image of the mid-point of $\alpha_{t_{1}}$ under $\rho_{t_{1}}$ lies in $(0,1)$ (so that $R(\alpha_{t_{1}})$ is symmetric with respect to the real line and $\rho_{t_{1}}(\alpha_{t_{1}})$ shrinks to the point $1\in\partial\mathbb{D}$), the map $\tilde{g}_{\mathbf{t},1}:=\rho_{t_{1}}\circ g_{\mathbf{t},1}\circ\rho_{t_{1}}^{-1}$ converges in Carathéodory topology (namely, uniformly on compact subsets) to the identity map in $\mathbb{D}$ as $t_{1}\to\infty$. If we Schwarz-reflect $g_{\mathbf{t},1}$ along $\partial\mathbb{D}\setminus g_{t_{1}}^{(1)}(\exp(\mathfrak{i}\theta_{2}))$, we see that the convergence also extends to the boundary, more precisely, we obtain that $\tilde{g}_{\mathbf{t},1}$ converges uniformly on all compact subsets of $\overline{\mathbb{D}}\setminus\{1\}$ (and the map is well-defined on every such compact subset for large enough $t_{1}$), so do the derivatives of $\tilde{g}_{\mathbf{t},1}$ with respect to $z$. Thus, we obtain that $R_{\mathbf{t}}\to 1$ almost surely, which completes the proof. ∎
@@ 3.4 Chordal SLE weighted by conformal radius
In this section, we show that the partition functions $\mathcal{Z}_{\alpha}$ correspond to the chordal SLE weighted by the conformal radius to the power $\alpha$.
For this, we first calculate the Laplace transform of the conformal radius of the complement of chordal SLE. Usually, chordal SLE is defined in the upper-half plane. It is more convenient here to describe it in the unit disc $\mathbb{D}$ via a change of coordinate. Fix $\kappa\in(0,8)$ and $\theta_{1}<\theta_{2}<\theta_{1}+2\pi$. Suppose $\gamma$ is chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$. We parameterize it by the capacity and define $g_{t},\xi_{t}$ accordingly as in Section 3.1. Denote by $T$ the first time $\gamma$ disconnects $\mathrm{e}^{\mathfrak{i}\theta_{2}}$ from the origin. A chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$, up to $T$, has the same law as radial $\operatorname{SLE}_{\kappa}(\kappa-6)$ starting from $\mathrm{e}^{\mathfrak{i}\theta_{1}}$ with force point $\mathrm{e}^{\mathfrak{i}\theta_{2}}$, up to the same time, see [41]. In other words, its driving function $\xi_{t}$ solves the following SDE:
@@
We first show that $\Phi(\kappa,\alpha;u)$ is finite as long as $\alpha<1-\kappa/8$. This is done in [36, Proof of Prop. 3.5]. For the reader’s convenience, we summarize its proof here.
@@
The disconnection time $T$ is the first time that $\theta_{t}$ hits 0 or $2\pi$. Suppose $f$ is an analytic function defined on $(0,1)$. Then $\mathrm{e}^{\alpha t}f\left(\left(\sin(\theta_{t}/4)\right)^{2}\right)$ is a local martingale if and only if $f$ satisfies (3.24). Since $f_{1},f_{2}$ are solutions to this ODE, the processes
@@ 3.5 Commuting SLEs without interchangeability
[…]
@@
[…]
@@ 4.1 Commutation relation when $\kappa=0$
We consider an interchangeable and locally commuting $2$-radial $\operatorname{SLE}_{0}$. Let $b_{1},b_{2}:S^{1}\times S^{1}\setminus\Delta\to\mathbb{R}$ be $C^{2}$ functions as in the condition (MARG). Then (2.4) and (2.6) imply that there exists $\mathcal{U}:\{(\theta_{1},\theta_{2})\in\mathbb{R}^{2}\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}$ and a constant $F$ such that
@@
$\begin{split}(\partial_{2}\mathcal{U})^{2}+2\cot(\theta_{12}/2)\partial_{1}\mathcal{U}-\frac{3}{\left(\sin(\theta_{12}/2)\right)^{2}}&=F,\\ (\partial_{1}\mathcal{U})^{2}+2\cot(\theta_{21}/2)\partial_{2}\mathcal{U}-\frac{3}{\left(\sin(\theta_{21}/2)\right)^{2}}&=F,\end{split}$ (4.1)
@@
Remark 4.2.
It follows immediately from the proof that, if one removes the interchangeability assumption from Proposition 4.1, then the only additional solutions $\mathcal{U}$ are of the from
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)+\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)-\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
for $\lambda>0$, up to an additive constant.
@@
The proof of Proposition 2.3 up to (2.11) holds verbatim when $\kappa=0$. By conformal invariance $b_{j}$ is invariant under rotation $(\theta_{1},\theta_{2})\mapsto(\theta_{1}+a,\theta_{2}+a)$. Thus, $\hat{b}_{j}:S^{1}\setminus\{0\}\to\mathbb{R}$ defined by $\hat{b}_{j}(\theta)=b_{j}(\theta,0)$, $j=1,2$, satisfies
$\partial_{1}b_{j}(\theta_{1},\theta_{2})=\partial_{1}\hat{b}_{j}^{\prime}(\theta_{12})=-\partial_{2}b_{j}(\theta_{1},\theta_{2}).$
Equations (2.10) and (2.11) give:
$(\cot(\theta/2)-\hat{b}_{2}(\theta))\hat{b}_{1}^{\prime}(\theta)-\frac{\hat{b}_{1}(\theta)}{2\sin^{2}(\theta/2)}+\frac{3\cot(\theta/2)}{2\sin^{2}(\theta/2)}=0$ (4.4)
$(\cot(\theta/2)+\hat{b}_{1}(\theta))\hat{b}_{2}^{\prime}(\theta)-\frac{\hat{b}_{2}(\theta)}{2\sin^{2}(\theta/2)}-\frac{3\cot(\theta/2)}{2\sin^{2}(\theta/2)}=0.$ (4.5)
Taking the difference of these equations yields
$\frac{\mathrm{d}}{\mathrm{d}\theta}((\cot(\theta/2)+\hat{b}_{1}(\theta))(\cot(\theta/2)-\hat{b}_{2}(\theta))+\frac{4\cot(\theta/2)}{\sin^{2}(\theta/2)}=0,$
so that
$(\cot(\theta/2)+\hat{b}_{1}(\theta))(\cot(\theta/2)-\hat{b}_{2}(\theta))=\frac{4}{\sin^{2}(\theta/2)}+C,$ (4.6)
for some $C\in\mathbb{R}$. Thus, multiplying (4.4) by $2\sin^{2}(\theta/2)(\hat{b}_{1}(\theta)+\cot(\theta/2))$ yields
$0=$ $(8+2C\sin^{2}(\theta/2))\hat{b}_{1}^{\prime}(\theta)-(\hat{b}_{1}(\theta)+\cot(\theta/2))(\hat{b}_{1}(\theta)-3\cot(\theta/2))$
$=$ $(8+2C\sin^{2}(\theta/2))f^{\prime}(\theta)-((f(\theta))^{2}+C+4)$
where $f$ is defined by $f(\theta)=\hat{b}_{1}(\theta)-\cot(\theta/2)$. Thus, $f$ solves the separable differential equation
$(8+2C\sin^{2}(\theta/2))f^{\prime}(\theta)=(f(\theta))^{2}+C+4.$ (4.7)
Similarly, we see using (4.5) that $g$ defined by $g(\theta)=\hat{b}_{2}(2\pi-\theta)-\cot(\theta/2)$ also solves (4.7) and by (4.6)
$(2\cot(\theta/2)+f(\theta))(2\cot(\theta/2)-g(2\pi-\theta))=\frac{4}{\sin^{2}(\theta/2)}+C.$ (4.8)
The solutions of (4.7) defined on $(0,2\pi)$ are studied in Lemma B.1.
Case 1: Suppose $C+4>0$. Then both $f$ and $g$ are of the form (B.2) for some constants $D=D_{f},D_{g}\in[-\pi/4,\pi/4].$ From (4.8) it follows that
$(C+4)=f(\pi)g(\pi)=(C+4)\tan(D_{f})\tan(D_{g})$
and thus we must have $D_{g}=D_{f}\in\{\pi/4,-\pi/4\}.$ Using the half-angle formula for the tangent function one can see that these two solutions can be rewritten as
$\begin{cases}f(\theta)=\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2)\\ g(\theta)=-\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2)\end{cases}$
$\begin{cases}f(\theta)=-\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2)\\ g(\theta)=\sqrt{(C+4)+4\cot^{2}(\theta/2)}-\cot(\theta/2).\end{cases}$
Case 2: If $C+4=0$ we have only the solutions
$\begin{cases}f(\theta)=0\\ g(\theta)=0\end{cases}\qquad\text{and}\qquad\begin{cases}f(\theta)=-4\cot(\theta/2)\\ g(\theta)=-4\cot(\theta/2).\end{cases}$
Indeed, if
[…]
for some $D\leq 0$ and $E\geq 0$ then solving (4.8) for $g$ yields
[…]
Therefore, we must have $D=E=0$ (otherwise $g$ is not well-defined on all of $(0,2\pi)$).
Case 3: For $C+4<0$ we have the solutions
$\begin{cases}f(\theta)=\sqrt{-C-4}\\ g(\theta)=\sqrt{-C-4}\end{cases}\quad\text{and}\quad\begin{cases}f(\theta)=-\sqrt{-C-4}\\ g(\theta)=-\sqrt{-C-4}.\end{cases}$
Here we have again used that $f(\pi)g(\pi)=-(C+4).$
From the solutions $(f,g)$ above we obtain corresponding solutions $(\hat{b}_{1},\hat{b}_{2})$ of (4.4) and (4.5) and $(b_{1},b_{2})$ of (2.10) and  (2.11) for $\kappa=0$. We note that these solutions satisfy $\partial_{1}b_{2}=\partial_{2}b_{1}$ (or equivalently $\hat{b}_{1}^{\prime}=-\hat{b}_{2}^{\prime}$, or $f^{\prime}(\theta)=g^{\prime}(2\pi-\theta)$) so there exists a function
$\mathcal{U}:\{(\theta_{1},\theta_{2})\,|\,\theta_{1}<\theta_{2}<\theta_{1}+2\pi\}\to\mathbb{R}$
such that $b_{1}=\partial_{1}\mathcal{U}$ and $b_{2}=\partial_{2}\mathcal{U}$. Up to an additive constant we have that
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)+\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2}):=-2\log\sin(\theta_{21}/2)-\int_{\theta_{21}}^{\pi}\sqrt{2\lambda+4\cot^{2}(u/2)}\mathrm{d}u,$
for $2\lambda=C+4>0$,
$\mathcal{U}(\theta_{1},\theta_{2})=-6\log\sin(\theta_{21}/2),$
and
$\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{\mu}(\theta_{1},\theta_{2})=2\log\sin(\theta_{21}/2)+\mu(\theta_{1}+\theta_{2})$
for $\mu=\pm\sqrt{-C-4}\in\mathbb{R}$. It is easily verified that each of these $\mathcal{U}$ satisfy (4.1) for some constant $F$. We note that $\mathcal{U}_{\lambda,\pm}$ does not satisfy the interchangeability condition but that the others do. ∎
@@
We note that when $\mathcal{U}(\theta_{1},\theta_{2})=\mathcal{U}_{0}(\theta_{1},\theta_{2})=2\log\sin(\theta_{21}/2)$, the corresponding commuting $\operatorname{SLE}_{0}$ is the two-sided radial $\operatorname{SLE}_{0}$ (with zero spiraling rate). This is also the geodesic pair in $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$ studied in [43, 29, 30], or in other words, the concatenation of $\eta^{(1)}$ and $\eta^{(2)}$ is the chord minimizing the chordal Loewner energy among all chords in $(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})$ passing through 0, hence can be viewed as the chordal $\operatorname{SLE}_{0}$ “conditioned” to pass through 0.